# FUSELOC: Fusing Global and Local Descriptors to Disambiguate 2D-3D Matching in Visual Localization

### v1

@@ IV-C Query time
where . This additional step significantly improves accuracy at the cost of increased memory usage (; see Table III). Finally, we search for the nearest neighbors of the combined descriptors in the codebook to obtain 2D-3D matching pairs.
V Experiments
We first provide implementation details in Section V-A. We evaluate our method on four popular datasets, which include outdoor scenes ranging from large buildings to city-level scales (Section V-B). We use the visual localization benchmark website111https://www.visuallocalization.net/benchmark/ to evaluate the performance of all methods. The weights for the global descriptor networks [8, 9, 11] and feature detectors [28, 27] are obtained off-the-shelf, without any re-training or fine-tuning on the benchmark datasets.
V-A Implementation details
In the experiments reported in Tables I and II, we set our only hyper-parameter . We note that is not the optimal parameter choice, leading to lower performance values than those that could be obtained by our method; we provide an ablation study for in Section VII-A. For each global descriptor method, we use the highest-performing variant as recommended by the authors. We use 16-bit floating-point precision to store the codebook descriptors, and the FAISS [30] library to facilitate nearest neighbor lookup using a GPU. Final poses are estimated using RANSAC provided by PoseLib [47].
V-B Datasets
Cambridge Landmarks [12] was recorded with a smartphone at five locations within the University of Cambridge, capturing a realistic urban environment with diverse lighting conditions and weather scenarios. Both training and testing sets were derived from multiple walking trajectories, introducing complexity to the localization task. The dataset contains ground-truth camera poses and 3D models as obtained by VisualSfM222http://ccwu.me/vsfm/index.html.
@@ V-B Datasets
| 0.25m/2∘ | 0.5m/5∘ | 5m/10∘ | 0.25m/2∘ | 0.5m/5∘ | 5m/10∘ | 0.25m/2∘ | 0.5m/5∘ | 5m/10∘ | Average (%) | ||
| Hierarchical methods | hLoc (SP+SG) [49, 50] | 83.4 | 93.4 | 99.7 | 52.0 | 87.2 | 96.1 | 92.9 | 94.5 | 95.6 | 88.3 |
| MegLoc [5] | 84.0 | 95.0 | 99.9 | 59.0 | 92.7 | 100.0 | - | - | - | - | |
| Image retrieval methods | MixVPR [8] | 0.0 | 0.4 | 27.4 | 6.6 | 23.7 | 79.8 | 9.2 | 28.4 | 96.0 | 30.2 |
| EigenPlaces [9] | 0.0 | 0.6 | 27.2 | 4.7 | 19.0 | 68.6 | 7.9 | 25.6 | 94.8 | 27.6 | |
| SALAD [11] | 0.0 | 0.5 | 27.8 | 6.3 | 22.8 | 96.6 | 7.2 | 23.0 | 96.6 | 31.2 | |
| CRICA [52] | 0.0 | 0.2 | 28.6 | 6.2 | 22.6 | 86.6 | 7.8 | 24.5 | 96.4 | 30.3 | |
| Structure-based methods (local descriptor only) | AS (SIFT) [2] | - | - | - | - | - | - | 63.0 | 69.9 | 78.5 | - |
| R2D2 [27] (vanilla) | 68.9 | 76.7 | 85.7 | 32.4 | 51.8 | 59.4 | 55.9 | 60.1 | 68.6 | 62.2 | |
| SIFT [23] (vanilla) | 56.0 | 60.0 | 65.6 | 24.9 | 39.2 | 43.6 | 34.6 | 38.6 | 45.4 | 45.3 | |
| SuperPoint [29] (vanilla) | 67.4 | 77.7 | 85.6 | 31.0 | 51.4 | 61.1 | 60.1 | 65.0 | 72.6 | 63.5 | |
| D2 [28] (vanilla) | 72.6 | 80.4 | 87.7 | 36.4 | 63.7 | 74.8 | 78.8 | 83.8 | 89.8 | 74.2 | |
| Structure-based methods (local and global descriptor) | D2 [28] + MixVPR [8] (light, ours) | 72.4 | 81.4 | 89.0 | 37.4 | 64.7 | 76.6 | 83.0 | 88.1 | 93.9 | 76.3 |
| D2 [28] + MixVPR [8] (heavy, ours) | 77.4 | 86.3 | 91.4 | 42.0 | 73.9 | 89.2 | 87.5 | 92.6 | 97.4 | 82.0 | |
| D2 [28] + EigenPlaces [9] (light, ours) | 76.9 | 86.4 | 93.2 | 36.6 | 63.8 | 76.1 | 85.2 | 90.5 | 95.8 | 78.3 | |
| D2 [28] + EigenPlaces [9] (heavy, ours) | 78.4 | 89.8 | 95.7 | 39.2 | 67.8 | 77.0 | 88.1 | 93.0 | 97.3 | 80.7 | |
| D2 [28] + SALAD [11] (light, ours) | 71.7 | 81.9 | 89.4 | 39.1 | 68.4 | 81.6 | 82.6 | 87.9 | 93.6 | 77.4 | |
| D2 [28] + SALAD [11] (heavy, ours) | 75.8 | 84.5 | 90.9 | 42.0 | 76.2 | 93.0 | 85.2 | 90.3 | 95.7 | 81.5 | |
| D2 [28] + CRICA [52] (light, ours) | 73.2 | 81.8 | 89.4 | 37.1 | 66.4 | 77.6 | 81.6 | 86.6 | 92.5 | 76.2 | |
| D2 [28] + CRICA [52] (heavy, ours) | 76.1 | 84.2 | 90.9 | 39.2 | 70.6 | 85.6 | 83.7 | 88.8 | 94.4 | 79.3 | |
@@
The Aachen Day-Night v1.1 Dataset [13] enhances the original Aachen dataset [15] with new sequences to construct a comprehensive 3D model of the historic inner city of Aachen, Germany, using COLMAP [46]. Training images were captured during daytime, while the test set includes nighttime images processed using HDR software to improve illumination.
RobotCar Seasons v2 [14] encompasses 20 million images collected over one year in a variety of weather conditions using an autonomous car equipped with six cameras, covering over 1000 km in Oxford, UK. For benchmarking, 49 different sub-models (each covering different locations) were reconstructed using high-quality images captured under overcast conditions. The test set includes images taken under a broader range of weather conditions. We evaluated algorithms against a global model containing the entire map rather than individual sub-models.
@@
VI Evaluation
VI-A Qualitative comparison
We qualitatively showcase the impact of our design on the Great Court sequence of the Cambridge Landmarks dataset [12]. This sequence features a large square dominated by overlapping building structures. To illustrate the perceptual aliasing problem in the codebook, we cluster all descriptors into 5 clusters with the K-means algorithm [31]. Figure 2 shows the point cloud for each cluster, illustrating how similar descriptors distribute across the map. When the codebook is trained using only local descriptors, we observe that 4 out of 5 clusters are spread across the whole map, confirming highly aliased descriptors throughout the map. Using our fused local+global descriptors, each cluster attends to different map sections due to the discriminating nature of the global descriptors. This significantly reduces the ambiguity of the search space, leading to a significant decrease in the number of false matches (see Figure 3).
VI-B Quantitative comparison
@@
Our method consistently outperforms the local-descriptor-only codebooks across all datasets (Tables I and II). For example, the average median translation error is reduced from cm to cm when fusing D2 with MixVPR on the Cambridge Landmarks [12] dataset (Table I). On larger maps (Aachen Day/Night v1.1 [13], RobotCar Seasons v2 [14] and Extended CMU seasons [15]), our method improves the percentage of successfully localized test images over the D2 local descriptor [28] from to on average, depending on the variant used (Table II).
@@
Our method further narrows the performance gap between direct matching algorithms [1, 2, 3] and hierarchical algorithms [4, 5], while retaining the appealing memory consumption of direct matching methods. On the Cambridge Landmarks dataset [12], we achieve an improvement of almost 1 cm (7.7%) in average median translation error over hLoc [4] while using only around 5% of the memory required (Table I). On larger maps with substantial perceptual aliasing, our method performs only worse compared to hLoc (Table II) while using only the memory footprint on average (Table III).
Note that this performance is achieved with which is not the optimal setting for our method. Section VII-A shows the optimal that renders our method competitive against hierarchical methods.
VI-C Memory requirements
@@
Table III measures the disk size of the required components for each method, including the codebook, the database images and their descriptors, and the mapping from database image pixels to point cloud coordinates. Since all methods under consideration require the point cloud coordinates and the weights of the deep networks, we exclude their memory footprints from the values reported in Table III. Depending on the dataset and whether our light or heavy variant is used, the storage requirements are roughly half of that of hLoc.
VII Ablation studies
VII-A Weights of local and global descriptors
@@
We conducted extensive experiments to examine the sensitivity of the parameter on two datasets: Aachen Day/Night v1.1 [13] and RobotCar Seasons v2 [14]. Using the heavy variant of our system, we tested ten different values ranging from to . Note that corresponds to the vanilla codebook, while is the value used for the results reported in Tables I and II.
@@
Figure 4 shows the percentage of successfully localized images for each value. We found that the optimal for SALAD [11] and CRICA [52] lies between and . This setting results in a codebook that significantly enhances performance compared to the default , achieving performance levels very close to hierarchical algorithms [4] (only and performance reduction on Aachen and RobotCar, respectively) while using less memory (Table III). Furthermore, we note that our system performs well for a wide range of , improving upon the vanilla system that only uses local descriptors for .
VII-B Global descriptor truncation
@@
gaussian Gaussian random projection [48],
random-0 random order with seed ,
first keep only the first entries,
center keep only the entries around the middle index,
last keep only the last entries.
Using the D2+SALAD variant, we again varied from to . Figure 5 shows that most truncation methods perform comparably, with random-0 performing slightly better than the other techniques.
VIII Conclusion
We introduce a simple technique to enhance 2D-3D search in direct matching visual localization. Through extensive evaluation on four real-world datasets, we demonstrate that our method significantly improves the performance of a brute-force baseline system with minimal memory overhead. Our ablation studies show that our approach performs comparably to hierarchical methods while using less memory, making our method particularly appealing for robotic systems with limited on-device memory.
The primary limitation of our method is its ability to disambiguate only non-co-visible points in the database map. We recommend future research to focus on identifying potential clues for resolving ambiguities among co-visible points. We hope our work continues to spark the community’s interest in 2D-3D matching systems because of their lower memory consumption and potential performance when combined with disambiguation techniques.

### v2

@@ 4.3 Query time
where . This additional step significantly improves accuracy at the cost of increased memory usage (by ; see Table 3).
Finally, we search for the nearest neighbors of the combined descriptors in the codebook to obtain 2D-3D matching pairs.
5 Experiments
We first provide implementation details and hardware settings in Section 5.1. We evaluate our method on four popular datasets, which include outdoor scenes ranging from large buildings to city-level scales (Section 5.2). We use the visual localization benchmark website111www.visuallocalization.net/benchmark/ for evaluation. The weights for the global descriptor networks [1, 7, 19] and feature detectors [15, 37] are obtained off-the-shelf, without any re-training or fine-tuning on the benchmark datasets.
5.1 Implementation details
We performed ablation experiments (see Section 7.1) to determine the optimal parameter for each global descriptor and reported their performance in Tables 1 and 2. We note that is determined only once during the codebook generation process and is adopted at query time. For each global descriptor method, we use the highest-performing variant as recommended by the authors. We use 16-bit floating-point precision to store the codebook descriptors, and the FAISS [20] library to facilitate nearest neighbor lookup using a GPU. Final poses are estimated using RANSAC provided by PoseLib [24]. Our experiments were performed using a Linux system with a single NVIDIA A100 GPU and 100 GB of RAM. Note that the codebook training process is memory intensive and can consume up to 100 GB. However, this requirement drops to less than 10 GB at query time (see Table 3).
5.2 Datasets
@@
| Memory requirement | 0.25m/2∘ | 0.5m/5∘ | 5m/10∘ | 0.25m/2∘ | 0.5m/5∘ | 5m/10∘ | 0.25m/2∘ | 0.5m/5∘ | 5m/10∘ | Average (%) | |
| hLoc (SP+SG) [38, 39] | 6.56 GB | 83.4 | 93.4 | 99.7 | 52.0 | 87.2 | 96.1 | 90.7 | 93.9 | 96.0 | 88.0 |
| GLACE [50] | 13 MB | 4.8 | 10.9 | 40.6 | - | - | - | - | - | - | - |
| AS (SIFT) [41] | - | - | - | - | - | - | - | 63.0 | 69.9 | 78.5 | - |
| R2D2 [37] | 0.75 GB | 68.9 | 76.7 | 85.7 | 32.4 | 51.8 | 59.4 | 55.9 | 60.1 | 68.6 | 62.2 |
| SIFT [28] | 0.75 GB | 56.0 | 60.0 | 65.6 | 24.9 | 39.2 | 43.6 | 34.6 | 38.6 | 45.4 | 45.3 |
| SuperPoint [14] | 1.5 GB | 67.4 | 77.7 | 85.6 | 31.0 | 51.4 | 61.1 | 60.1 | 65.0 | 72.6 | 63.5 |
| D2 [15] | 3 GB | 72.6 | 80.4 | 87.7 | 36.4 | 63.7 | 74.8 | 78.8 | 83. | 89.8 | 74.2 |
| Ours (light, D2+SALAD, ) [15, 19] | 3 GB | 77.2 | 87.2 | 93.0 | 42.7 | 78.2 | 94.8 | 85.9 | 91.1 | 96.7 | 83.0 |
| Ours (light, D2+MegaLoc, ) [15, 6] | 3 GB | 74.8 | 84.7 | 91.2 | 39.2 | 67.7 | 80.6 | 83.6 | 88.8 | 94.3 | 78.3 |
| Ours (light, D2+MixVPR, ) [15, 1] | 3 GB | 72.4 | 81.4 | 89.0 | 37.4 | 64.7 | 76.6 | 83.0 | 88.1 | 93.9 | 76.3 |
[…]
| Ours (heavy, D2+MixVPR, ) [15, 1] | 3.13 GB | 77.4 | 86.3 | 91.4 | 42.0 | 73.9 | 89.2 | 87.5 | 92.6 | 97.4 | 82.0 |
@@
| hloc [38] | ours (heavy) | ours (light) | |
| Aachen | 0.19 | 0.30 | 0.30 |
| RobotCar | 0.32 | 0.56 | 0.57 |
| CMU | 0.40 | 0.58 | 0.69 |
| Average | 0.3 | 0.48 | 0.52 |
The Cambridge Landmarks Dataset [21] was collected at the University of Cambridge using a Google LG Nexus 5 smartphone in large-scale outdoor urban environments. HD videos were recorded at 2 Hz to extract RGB frames. The dataset features typical urban challenges such as moving pedestrians and vehicles, and captures different lighting and weather conditions over time. Training and test data were obtained from disjoint walking trajectories, increasing the difficulty of the localization task. Ground-truth camera poses and 3D models were generated using VisualSfM [53] and are provided with the dataset.
The Aachen Day-Night v1.1 Dataset [56] enhances the original Aachen dataset [44] with new sequences to construct a comprehensive 3D model of the historic inner city of Aachen, Germany, using COLMAP [45]. Training images were captured during the daytime, while the test set includes nighttime images processed using HDR software to improve illumination.
RobotCar Seasons v2 [31] encompasses 20 million images collected over a one year period in a variety of weather conditions using an autonomous car equipped with six cameras, covering over 1000 km in Oxford, UK. For benchmarking, 49 different sub-models (each covering different locations) were reconstructed using high-quality images captured under overcast conditions. The test set includes images taken under a broader range of weather conditions. We evaluated algorithms against a global model containing the entire map rather than individual sub-models.
@@
6 Evaluation
6.1 Quantitative comparison
@@
Our method consistently outperforms the local-descriptor-only codebooks across all datasets (Tables 1 and 2). For example, the average median translation error is reduced from cm to cm when fusing D2 with MixVPR on the Cambridge Landmarks [22] dataset (Table 1). On larger maps (Aachen Day/Night v1.1 [56], RobotCar Seasons v2 [31] and Extended CMU seasons [44]), our method improves the percentage of successfully localized test images over the D2 local descriptor [15] from to on average, depending on the variant used (Table 2). We refer the reader to Figure 3 for a complete overview of evaluation results averaged over these three large-scale datasets for different combinations of global and local descriptors.
@@
Our method further narrows the performance gap between direct matching algorithms [26, 41, 40] and hierarchical algorithms [38, 35], while retaining the appealing computational complexity of direct matching methods. On the Cambridge Landmarks dataset [22], we achieve an improvement of almost 1 cm () in average median translation error over hLoc [38] while using only around of the memory required (Table 1). On larger maps with substantial perceptual aliasing, our method performs only worse than hLoc (Table 2) while using only of the memory footprint (Table 3) and being times faster on average (Table 4).
6.2 Computational complexity
@@
Table 3 measures the memory requirements of the required components for each method, including the codebook, the database images and their descriptors, and the mapping from database image pixels to point cloud coordinates. Since all methods under consideration require the point cloud coordinates and the weights of the deep networks, we exclude their memory footprints from the values reported in Table 3. Depending on the dataset and whether our light or heavy variant is used, the memory requirements are roughly half of that of hLoc [38].
We report our frame rate in Table 4. By performing direct matching against 3D points, we eliminate the costly feature detection step for database images. While caching these database features could accelerate hierarchical algorithms, it would demand approximately 30-50 times more memory. For example, storing only the D2 [15] local features consumed 158.4 GB for the Aachen dataset and 223.3 GB for the RobotCar dataset. Given the memory settings outlined in Table 3, our method achieves a speedup of approximately 1.6x compared to hLoc [38].
7 Ablation studies
7.1 Hyperparameter
@@
We conducted extensive experiments to examine the sensitivity of the parameter on two datasets: Aachen Day/Night v1.1 [56] and RobotCar Seasons v2 [31]. Using the heavy variant of our system, we tested ten different values ranging from to . Note that corresponds to the vanilla codebook.
@@
Figure 4 shows the percentage of successfully localized images for each value. We found that the optimal for SALAD [19] and CRICA [30] lies between and . This setting results in a codebook that significantly enhances performance compared to the default , achieving performance levels very close to hierarchical algorithms [38] (only and performance reduction on Aachen and RobotCar, respectively) while using less memory (Table 3) and being times faster (Table 4). Furthermore, we note that our system performs well for a wide range of , consistently improving upon the vanilla system that only uses local descriptors for .
7.2 Global descriptor truncation
@@
Using the D2+SALAD variant, we again varied from to . Figure 5 shows that most truncation methods perform comparably, with random-0 and pca performing slightly better than the other techniques. Therefore, we use random-0 for all of our experiments due to its simplicity.
8 Conclusion
We introduce a simple technique to enhance 2D-3D search in direct matching visual localization. Through extensive evaluation on four real-world datasets, we demonstrate that our method significantly improves the performance of a brute-force baseline system with minimal memory overhead. Our ablation studies show that our approach performs comparably to hierarchical methods while using less memory and running times faster, making our method particularly appealing for robotic systems with limited onboard memory.
The primary limitation of our method is that it can only disambiguate non-co-visible points in the database map. We recommend that future research focuses on identifying potential clues for resolving ambiguities among co-visible points. We hope our work will continue to spark the community’s interest in 2D-3D matching systems due to their lower memory consumption and high potential performance when combined with disambiguation techniques.
Acknowledgements
This research was partially supported by funding from ARC Laureate Fellowship FL210100156 to MM, ARC DECRA Fellowship DE240100149 to TF, and the QUT Centre for Robotics.